Assume that there are only two countries: Eastland and Westland both producing cups and plates. These two countries have identical demand, production technology, and resource endowment. Each country has 6,000 labor hours available of which it allocates equally to each industry. In one hour, each country can produce 5 cups and 3 plates. The labor productivity for plates does not change with the quantity of output. On the other hand, when Eastland produces 30,000 or more cups, its productivity increases to 6 cups per hour.

A) What is the total world consumption after trade?

B) What is the range of terms of trade for 1 cup?

C) What is the range of terms of trade for 1 plate?
Answer #1

When Eastland produces 30,000 or more cups, its productivity increases to 6 cups per hour, therefore Eastland shall use all its labour hours in producing cups. Therefore, total number of cups produced by Eastland = 6000*6 = 36000

Westland will therefore, utilize its entire labor hours in producing plates, total number of plates produced by Westland = 6000*3 = 18000

A) Total world consumption after trade = 36000 cups and 18000 plates

B) Terms of trade for one cup: (3/6, 3/5) plates
Terms of trade will be upper bound by the fact that both that if East were to demand more than 3/5 of the plates for one cup, Westland will be better off producing it themselves.Conversely they will not ask for lower than 3/6, as in that case, they will be better off producing themselves.

C) Terms of trade for one plate: (5/3, 6/3)
Similar to the explanation above as the terms will inverse in the case of plate